You left your house driving at 40 mph and your car broke down. You had your bike so you rode all the way home at 10 mph. The entire trip took you four hours. How far from your house did you break down?
Please break this down as simple as possible. I want to get it to "click" in my head so I can get it. Thanks!!!

Answer by htmentor(1343)

Let d = the distance travelled before breaking down.
For the outbound trip in the car, the travel time = d/40 [time = distance/speed]
For the return trip on the bike, the travel time = d/10 [same distance travelled]
The total time = 4 hr, so we can write
d/40 + d/10 = 4
Solve for d:
(40d + 10d)/400 = 4
50d = 1600
d = 32
So you travelled 32 miles before breaking down